OpenStudy (anonymous):

How do you find a b and c for a=-b+3 a+2c=10 -b+c=3

jimthompson5910 (jim_thompson5910):

Solve -b+c=3 for c to get c = b+3

jimthompson5910 (jim_thompson5910):

Now plug a = -b+3 into the second equation to get a+2c=10 (-b+3)+2c=10 -b+3+2c=10 Then plug in c = b+3 to get -b+3+2c=10 -b+3+2(b+3)=10 Solve this last equation for b. Tell me what you get.

OpenStudy (anonymous):

What do you plug c=b+3 into?

jimthompson5910 (jim_thompson5910):

into -b+3+2c=10

jimthompson5910 (jim_thompson5910):

notice how I replaced 'c' with b+3

OpenStudy (anonymous):

thanks

OpenStudy (anonymous):

b=1?

jimthompson5910 (jim_thompson5910):

yes good

jimthompson5910 (jim_thompson5910):

you can now use this to find 'a' and 'c'

OpenStudy (anonymous):

a=2

OpenStudy (anonymous):

c=4

jimthompson5910 (jim_thompson5910):

good

jimthompson5910 (jim_thompson5910):

all 3 are correct

OpenStudy (anonymous):

thanks!

jimthompson5910 (jim_thompson5910):

yw